Circular No. 6053

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET MACHHOLZ (1994o) Donald E. Machholz reports his visual discovery of a comet. The fol-

lowing observations are available:

1994 UT  α_{2000}  δ_{2000}  m_{1} Observer
Aug. 13.4215 4^{h}12.^{m}9  +62^{o}48′  10 Machholz
13.80190 4 20 30.42 +62 37 18.1  Kojima
13.80304 4 20 31.66 +62 37 18.1  _{”}
13.80431 4 20 33.44 +62 37 18.6  _{”}
13.80535 4 20 35.32 +62 37 11.6  _{”}

D. E. Machholz (Colfax, CA). 0.25-m reflector. Comet diffuse with little

condensation; coma diameter perhaps 3′–4′. T. Kojima (YGCO Chiyoda Observatory). 0.25-m reflector + CCD. Poor

conditions in twilight. Comet diffuse with condensation. Communicated by S. Nakano, Sumoto, Japan.

JUPITER AND COMET 1993e G. Bjoraker, Goddard Space Flight Center, NASA; and T. Herter, S.

Stolovy, G. Gull, and B. Pirger, Cornell University, write: “We have detected cometary water in the fireball of fragments G and K immediately after the plume became visible to observers on Earth. The observations were performed using the Kuiper Echelle Grating Spectrometer (spectral resolving power 9000; 5′′ slit) on the Kuiper Airborne Observatory. The G fireball was observed at ∼ July 18.323 UT; the K fragment fireball was observed on July 19.444. Three H_{2}O lines were observed in emission for both the G and K fireballs at the following wavelengths: 7.7136, 7.7118, and 7.7090 μm. Numerous jovian ^{12}CH_{4} and ^{13}CH_{4} lines were observed in emission across our bandpass (7.671–7.722 μm). This serves as a thermometer of Jupiter’s stratosphere, sounding the 1-microbar to 10-millibar levels. The radiance in the jovian CH_{4} lines increased by a factor of ∼ 25– 30 between the pre-crash level and the peak of the fireball. The CH_{4} lines were observed to decay over the next 2 hr. The three H_{2}O lines have a distinctively different time behavior: they essentially disappear over a 30-min period. Very high temperatures (T > 500 K) are required to observe these H_{2}O transitions. The observing geometry, high temperature, and short time scale for the disappearance of these H_{2}O lines favor a cometary (rather than jovian) source. This, in turn, would confirm that P/Shoemaker-Levy 9 was a comet and not an asteroid.”

1994 August 13 Daniel W. E. Green